OpenStudy (anonymous):

Apply Euler's method with the step h = 1/2 to find an approximation of y(1) where y is the solution of the initial value problem y'=1+y y(0)=0

OpenStudy (anonymous):

i get y = 7/2

OpenStudy (anonymous):

y'=1+y , y(0)=0 y1=y0+(1+y0)h, y1=0 +(1+0)(0.05) =0.5

OpenStudy (anonymous):

the exact value y(1)=4.436563657